# Evaluating diversion and treatment policies for opioid use disorder

### v3

@@ Abstract
The United States (US) opioid crisis contributed to 81,806 fatalities in 2022. It has strained hospitals, treatment facilities, and law enforcement agencies due to the enormous resources and procedures needed to respond to the crisis. As a result, many individuals who use opioids never receive or finish the treatment they need and instead have many interactions with hospitals or the criminal justice system. This paper introduces a discrete event simulation model that evaluates three opioid use disorder treatment policies: arrest diversion, re-entry case management, and overdose diversion. Publicly available data from 2011 to 2019 in Dane County, Wisconsin, was used to forecast opioid-related outcomes through 2032. Through analyzing a variety of policy-mix implementations, the study offers a versatile framework for evaluating policies at various implementation levels. The results demonstrate that treatment policies that create new pathways and programming by utilizing treatment services and successfully divert at least 20% of eligible individuals can lead to more opioid-resilient communities. The benefits increase when more policies are enacted and/or offered to more individuals, with the largest impact from overdose diversion, followed by re-entry case management, and the smallest impact from arrest diversion. The cumulative total reduction in societal costs from 2017 to 2032 ranges from \$68 M (USD) to \$1.06 B (USD), excluding implementation costs of policies. To reverse the opioid crisis within a community, treatment policies may need to be combined with other strategies, such as harm reduction, supply reduction, and use prevention.
@@ 1 Introduction
There were 81,806 opioid-related deaths in 2022 in the US, a significant increase from 10,000 deaths in 1999 (National Center for Health Statistics,, 2022). The rapid increase in opioid-related deaths has become known as the “Opioid Crisis.” It is characterized by three waves: prescription opioids in the 1990s, heroin in the 2010s, and synthetic opioids since 2013 (Pardo et al.,, 2019). Most of the literature in industrial engineering focuses on limiting the available supply of opioids through stricter laws and network interdiction modeling and decisions (Anzoom et al.,, 2021). However, the opioid crisis requires a larger array of interventions (Pardo et al.,, 2019).
@@
SUD is defined as the uncontrollable chronic use of a substance (i.e., legal or illegal drugs) that causes clinically significant distress or impairment. opioid use disorder (OUD) is a type of substance use disorder (SUD) where the substance used is an opioid (e.g., tramadol, heroin, oxycodone, fentanyl). Not all individuals who use opioids are classified as having OUD and, instead, may use opioids occasionally or have regular controlled use that does not affect their well-being, function, or social relationships. However, all individuals who use opioids are at risk of opioid-related death and potential arrest in the majority of the US. In fact, individuals who use illicit drugs often have many interactions with police and the criminal justice system (CJS). These individuals are more likely to be arrested for various crimes related to their drug use, such as drug possession or sale, driving under the influence of drugs, burglary, and prostitution (National Institute on Drug Abuse,, 2018), which creates a cycle and often leads to time in prison. This cycle is costly for individuals and society (The Council of Economic Advisers,, 2019). It follows that illicit drug use heavily impacts the CJS. The US Department of Justice estimates that more than half of male and two-thirds of female state prisoners and sentenced jail inmates have a SUD, with 16.6% and 18.9% of state prisoners and sentenced jail inmates, respectfully, using heroin or other opiates (Bronson et al.,, 2017). Following release from prison, individuals are estimated to be between seven to twelve times more likely than the general population to die of an overdose (Binswanger et al.,, 2007; Bird and Hutchinson,, 2003).
@@
Several initiatives have aimed to break the cycle of recidivism and OUD through diversion programming, which redirects individuals to treatment and counseling instead of the CJS. One diversion program is arrest diversion, also known as officer intervention, where officers refer individuals suspected of drug-related crimes to mental health treatment, behavioral health counseling, or case management (COSSAP, 2020). Arrest diversion programs have shown promising effects on recidivism outcomes (PAARI, 2019; 2021; PAARI) but face various barriers, making it difficult to evaluate whether the program was executed as planned (Joudrey et al.,, 2021). Similarly, after someone serves their sentence and re-enters society re-entry case management can connect individuals to relevant resources, such as OUD treatment via a case manager or parole officer. Another treatment policy is overdose diversion, which works similarly to re-entry case management except for the medical system.
@@ 2.1 Literature Review
Some argue for applying mathematical modeling and simulation to these treatment policies using available data before implementation (Bansback et al.,, 2021). Simulation has proven useful for planning, implementation, and preparedness for various public sector issues, including public health crises. For example, during the COVID-19 pandemic, simulation informed policy decisions such as minimum staffing needs, social distancing requirements, and delivery of testing and vaccines (Currie et al.,, 2020; Cao and Liu,, 2022). Additionally, simulation has been applied to various applications in the healthcare system, from disease screening to health and care system operations (Zhang,, 2018). The need for a simulation modeling approach within the context of the opioid crisis is highlighted in literature reviews (e.g., Sharareh et al., (2019), Hoang et al., (2016)). Specifically, Sharareh et al., (2019) review opioid-related mathematical and conceptual models, concluding that more mathematical and analytical modeling is needed to tackle the complexity of the opioid crisis. Many simulations seek to evaluate if a specific program or treatment will reduce adverse outcomes (e.g., crime, fatal overdoses) and evaluate the program’s cost-effectiveness. Literature reviews by Beaulieu et al., (2021) and Barbosa et al., (2020) find a lack of simulations that model the trade-offs of implementing various interventions. We propose using a DES to test multiple OUD treatment policies. We chose DES as the type of simulation due to the discrete nature of an individual’s experience (e.g., an overdose, starting treatment, an arrest). In contrast to the existing papers in the literature, this paper models at the individual level while testing the effects of separate policies. We also follow the recommendations for opioid-related simulations by Cerdá et al., (2021), who conduct a literature review of parameters used in opioid-related simulations and conclude that many simulations do not meet standard transparency and reproducibility criteria.
@@ 2.2 Simulation Model
Individuals enter the simulation via distribution (1), representing an individual's first use or initiation of opioids. Upon entering the simulation, the individual is classified as actively using opioids for a period of time before the next event occurs. At the time of an individual’s first start opioid use, seven event times are generated that follow distributions (2)-(8). The seven possible next events are natural death, opioid-related death, opioid-related arrest, non-opioid-related arrest, opioid-related hospital encounter, stop opioid use, and the start of OUD treatment. The event with the earliest time is scheduled, and the others are discarded, except for the natural death time and non-opioid-related arrest. Note, we model an active use state rather than an OUD state. Therefore, individuals in the active use state could have anywhere from prescription use to casual misuse to severe OUD. No matter an individual’s dependence on opioids, they are still at risk of opioid-related death, having an opioid-related hospital encounter, and entering the CJS for opioid-related arrests such as drug paraphernalia, illicit sale, and possession. For OUD treatments, individuals who are referred to treatment or start treatment may or may not have OUD. This is realistic since individuals who arrive at a hospital or have an opioid-related arrest could have varying levels of use and treatment needs.
@@ 2.2.2 Model Outputs
The main outcomes of the model are the number of opioid- and non-opioid-related deaths, opioid- and non-opioid-related diverted and non-diverted arrests, OUD treatments, and opioid-related hospital encounters in a given year. The model also estimates the amount of time in and the number of individuals in each state at a given time, which can be used to predict demand for that given resource. Demand for hospital beds, CJS usage, and OUD treatment are not typically publicly available or known and are critical to understanding potential policy impacts on existing systems. Additionally, we estimate the total yearly societal costs of the opioid crisis for each OUD treatment policy scenario, ignoring the initial implementation costs. Our analysis compares each scenario’s cost difference and difference in outcomes against the base model to provide an additional understanding of the trade-offs involved in referring more individuals to OUD treatment.
@@ 2.3 Case study: Dane County, Wisconsin
The DES model is evaluated using a case study based in Dane County, WI. Dane County is in south-central Wisconsin and contains the metropolitan area of Madison. Only persons aged 12+ are included. Dane County had an estimated 12+ population of 473,436 in July 2019, making it a mid-size county in the U.S. (U.S. Census Bureau,, 2019). Dane County is a natural choice for a case study since it implemented an arrest diversion policy named the Madison Addiction Recovery Initiative (MARI) that began in August of 2017 (Madison Police Department,, 2020) as well as an overdose diversion policy, named ED2Recovery, in 2017 (Wisconsin Voices For Recovery,, 2021). This study addresses analytic questions that the MARI program could not directly study. For example, the MARI team only had data on MARI participants and could not address broader Dane County outcomes (White et al.,, 2021; Nyland et al.,, 2024). Additionally, since the program was not implemented in isolation, it is difficult to assess Dane County-level trends due to MARI or ED2Recovery. Therefore, the DES simulation model provides insight into the effects of the three different OUD treatment programs.
Table 1 reports model inputs used in the case study. Table 1 includes distributions, model inputs, and data sources used. The distributions include: exponential (Exp) with parameter $\lambda$, triangular (Tri) with parameters minimum, median, and maximum, and log-normal (LN) with parameters $\mu$, $\sigma$, and a location parameter of 0 unless otherwise stated. A LN distribution is used to estimate times in OUD treatment, CJS, and the hospital, since individuals are most likely to spend a short amount of time in jail, OUD treatment, and the hospital with a low likelihood of remaining in each of these states for an extended period of time. Due to the “new wave” of opioid deaths starting in 2019 caused by the introduction of fentanyl in the drug supply, we split arc (2) into pre-2019 (i.e., through the end of 2018) and post-2019 (i.e., 2019 and on) to capture the increased deadliness of opioids due to fentanyl (National Center for Health Statistics,, 2022).
@@
Dist. Parameter Distribution for Base Model Data Source
NA Age (years) at opioid initiation LN(2.88, 0.56) with Loc = 12 Lewer et al., (2020)
NA Age (years) in starting population LN(3.10, 0.46) with Loc = 12 Lewer et al., (2020)
NA Starting population size Tri(27299, 34224, 43261) SAMHSA (2019)
NA Starting state Multinomial, See Section 2.4 Expert Opinion
(1) Time of next arrival (days) Exp(10.87) Multiple years 2015-2019 SAMHSA (2019)
@@
(2) pre-2019 Time in active state given next event is an opioid-related death before 2019 (days) Right-skewed non-parametric version of LN(17.60, 4.19) National Center for Health Statistics, (2022) 2016-2018
@@
(2) post-2019 Time in active state given next event is an opioid-related death 2019 and later (days) Right-skewed non-parametric version of LN(17.13, 4.14) National Center for Health Statistics, (2022) 2019-2020
@@
(3) Time in active state given next event is a hospital encounter (days) Right-skewed non-parametric version of LN(9.07, 3.01) Wisconsin Department of Health Services, 2017b 2018-2019
@@
(4) Time in active state given next event is an opioid-related arrest (days) Right-skewed non-parametric version of LN(10.04, 2.35) Wisconsin Department of Justice, (2016) 2016-2019
@@
(5) Time in active state given next event is OUD treatment (days) Right-skewed non-parametric version of LN(7.48, 1.03) Wisconsin Department of Health Services, (2019) 2016-2017
@@
(6) Time in active state given next event is inactive use (days) Right-skewed non-parametric version of LN(4.82, 2.20) Rivera et al., (2018); Bauer, (2019)
@@
(7) Time until individual's non-opioid related death (days) Empirical, See Section A.5.7 Arias et al., (2013)
@@
(8) Time until individual's non-opioid related arrest (days) Right-skewed non-parametric version of LN(7.88, 2.38) Wisconsin Department of Justice, (2016) 2016-2019
(A) Hospital service time (days) LN(0.82, 0.48) Singh and Cleveland, (2020); Reinert et al., (2019)
(B) CJS service time (days) LN(2.16, 1.47) Zhang et al., (2022)
@@
(C) OUD treatment service time (days) LN(4.78, 1.18) Division of Care and Teatment Services, (2018)
@@
(D) Inactive service time after release from CJS (days) LN(3.29, 1.61) Bukten et al., (2017); Kinlock et al., (2008)
(E) Inactive service time after OUD treatment (days) LN(4.52, 1.09) Nunes et al., (2018)
@@
(F) Inactive service time after hospital or ED discharge (days) LN(1.95, 1.40) Chutuape et al., (2001)
@@
(G) Inactive service time after active state (days) LN(6.29, 2.51) Calibrated
$p_{D}$ of Hosp. encounters leading to death Probability(0.0218) Singh and Cleveland, (2020)
$p_{A}$ of Hosp. encounters leading to arrest Probability(0.01) Expert Opinion
$p_{AD}$ of Arrests leading to treatment Probability(0) Not Applicable
$p_{OD}$ of Hosp. encounters leading to treatment Probability(0.2227) Singh and Cleveland, (2020)
$p_{CM}$ of CJS re-entries leading to treatment Probability(0) Not Applicable
@@
Table 2 reports the 2017 USD average cost per event for opioid-related-death, opioid-related-arrest, start OUD treatment, hospital encounter, active opioid use at year-end, and inactive opioid use at year-end. We obtain the 2017 USD cost of a single opioid-related death of \$11,548,462 from Luo et al., (2021). We estimate the cost of an individual starting OUD treatment in Wisconsin to be \$$\$1,812$. The per-event cost of starting OUD treatment was estimated by dividing the total estimated Wisconsin burden in 2017 due to OUD treatment (i.e., \$$\$59.8$ million) from Luo et al., (2021) by the total number of 2017 OUD treatment episodes in the state of Wisconsin (i.e., $33,005$) obtained from Wisconsin Department of Health Services, (2019). We estimated the cost of a hospital encounter to be \$$\$20,077$. Similarly, the per-event cost of a hospital encounter was estimated by dividing the total estimated Wisconsin burden in 2017 due to health care (i.e., \$$\$529.4$ million) from Luo et al., (2021) by the total number of 2017 hospital encounters in the state of Wisconsin (i.e., $26,369$) from Wisconsin Department of Health Services, 2017b . We use the healthcare cost, estimated by Luo et al., (2021), as a surrogate cost of hospital encounters since hospital encounter-related costs dominate the total healthcare costs (Florence et al.,, 2021).
@@
Start OUD treatment \$1,812
Hospital encounter \$20,077

### v4

@@ Abstract
The United States (US) opioid crisis contributed to 81,806 fatalities in 2022. It has strained hospitals, treatment facilities, and law enforcement agencies due to the enormous resources and procedures needed to respond to the crisis. As a result, many individuals who use opioids never receive or finish the treatment they need and instead have many interactions with hospitals or the criminal justice system. This paper introduces a discrete event simulation model that evaluates three opioid use disorder treatment policies: arrest diversion, re-entry case management, and overdose diversion. Publicly available data from 2011 to 2019 in Dane County, Wisconsin, was used to forecast opioid-related outcomes through 2032. Through analyzing a variety of policy-mix implementations, the study offers a versatile framework for evaluating policies at various implementation levels. The results demonstrate that treatment policies that create new pathways and programming by utilizing treatment services and successfully divert at least 20% of eligible individuals can lead to more opioid-resilient communities. The benefits increase when more policies are enacted and/or offered to more individuals, with the largest impact from overdose diversion, followed by re-entry case management, and the smallest impact from arrest diversion. The statistically significant 10-year cumulative total reduction in societal costs from 2023 through 2032 ranges from \$39 M (USD) to \$584 M (USD), excluding implementation costs of policies. To reverse the opioid crisis within a community, treatment policies may need to be combined with other strategies, such as harm reduction, supply reduction, and use prevention.
Disclaimer: This is the accepted manuscript version of the article published in IISE Transactions on Healthcare Systems Engineering. This manuscript may differ slightly from the final version, which is available at https://doi.org/10.1080/24725579.2025.2523273.
@@ 1 Introduction
There were 81,806 opioid-related deaths in 2022 in the US, a significant increase from 10,000 deaths in 1999 (National Center for Health Statistics,, 2022). The rapid increase in opioid-related deaths has become known as the “Opioid Crisis.” It is characterized by three waves: prescription opioids in the 1990s, heroin in the 2010s, and synthetic opioids since 2013 (Pardo et al.,, 2019). Most of the literature in industrial engineering focuses on limiting the available supply of opioids through stricter laws and network interdiction modeling and decisions (Anzoom et al.,, 2021). However, the opioid crisis requires a wider array of interventions (Pardo et al.,, 2019).
@@
[…]
@@
Several initiatives have aimed to break the cycle of recidivism and OUD through diversion programming, which redirects individuals to treatment and counseling instead of the CJS. One diversion program is arrest diversion, also known as officer intervention, where officers refer individuals suspected of drug-related crimes to mental health treatment, behavioral health counseling, or case management (COSSAP, 2020). Arrest diversion programs have shown promising effects on recidivism outcomes (PAARI, 2019; 2021; PAARI) but face various barriers, making it difficult to evaluate whether the program was executed as planned (Joudrey et al.,, 2021). Similarly, after someone serves their sentence and re-enters society, re-entry case management can connect individuals to relevant resources, such as OUD treatment via a case manager or parole officer. Another treatment policy is overdose diversion, which works similarly to re-entry case management except for the medical system.
@@ 2.1 Literature Review
Some argue for applying mathematical modeling and simulation to these treatment policies using available data to predict policy performance before implementation (Bansback et al.,, 2021). Simulation has proven useful for planning, implementation, and preparedness for various public sector issues, including public health crises. For example, during the COVID-19 pandemic, simulation informed policy decisions such as minimum staffing needs, social distancing requirements, and delivery of testing and vaccines (Currie et al.,, 2020; Cao and Liu,, 2022). Additionally, simulation has been applied to various applications in the healthcare system, from disease screening to health and care system operations (Zhang,, 2018). The need for a simulation modeling approach within the context of the opioid crisis is highlighted in literature reviews (e.g., Sharareh et al., (2019), Hoang et al., (2016)). Specifically, Sharareh et al., (2019) review opioid-related mathematical and conceptual models, concluding that more mathematical and analytical modeling is needed to tackle the complexity of the opioid crisis. Many simulations seek to evaluate if a specific program or treatment will reduce adverse outcomes (e.g., crime, fatal overdoses) and evaluate the program’s cost-effectiveness. Literature reviews by Beaulieu et al., (2021) and Barbosa et al., (2020) find a lack of simulations that model the trade-offs of implementing various interventions. We propose using a DES to test multiple OUD treatment policies. We chose DES as the type of simulation due to the discrete nature of an individual’s experience (e.g., an overdose, starting treatment, an arrest). In contrast to the existing papers in the literature, this paper models at the individual level while testing the effects of separate policies. We also follow the recommendations for opioid-related simulations by Cerdá et al., (2021), who conduct a literature review of parameters used in opioid-related simulations and conclude that many simulations do not meet standard transparency and reproducibility criteria.
@@ 2.2 Simulation Model
Individuals enter the simulation via distribution (1), representing an individual's first use or initiation of opioids. Upon entering the simulation, the individual is classified as actively using opioids for a period of time before the next event occurs. At the time of an individual’s first start opioid use, seven event times are generated that follow distributions (2)-(8). The seven possible next events are natural death, opioid-related death, opioid-related arrest, non-opioid-related arrest, opioid-related hospital encounter, stop opioid use, and the start of OUD treatment. The event with the earliest time is scheduled, and the others are discarded, except for the natural death time and non-opioid-related arrest. Note, we model an active use state rather than an OUD state. Therefore, individuals in the active use state could have anything from prescription use to casual misuse to severe OUD. No matter an individual’s dependence on opioids, they are still at risk of opioid-related death, having an opioid-related hospital encounter, and entering the CJS for opioid-related arrests such as drug paraphernalia, illicit sale, and possession. For OUD treatments, individuals who are referred to treatment or start treatment may or may not have OUD. This is realistic since individuals who arrive at a hospital or have an opioid-related arrest could have varying levels of use and treatment needs.
@@ 2.2.2 Model Outputs
The main outcomes of the model are the number of opioid- and non-opioid-related deaths, opioid- and non-opioid-related non-diverted arrests, OUD treatments, and opioid-related hospital encounters in a given year. The model also estimates the amount of time in and the number of individuals in each state at a given time, which can be used to predict demand for that given resource. Demand for hospital beds, CJS usage, and OUD treatment are not typically publicly available or known and are critical to understanding potential policy impacts on existing systems. Additionally, we estimate the total yearly societal costs of the opioid crisis for each OUD treatment policy scenario, ignoring the initial implementation costs. Our main analysis reports each scenario’s 10-year cumulative outcomes starting in 2023 through the end of 2032. We then compare the difference in outcomes against the base model to provide additional understanding of the trade-offs involved in referring more individuals to OUD treatment.
@@ 2.3 Case study: Dane County, Wisconsin
The DES model is evaluated using a case study based in Dane County, which is in south-central Wisconsin and contains the metropolitan area of Madison. Only persons aged 12+ are included. Dane County had an estimated 12+ population of 473,436 in July 2019, making it a mid-size county in the U.S. (U.S. Census Bureau,, 2019). Dane County is a natural choice for a case study since it implemented an arrest diversion policy named the Madison Addiction Recovery Initiative (MARI) that began in August of 2017 (Madison Police Department,, 2020) as well as an overdose diversion policy, named ED2Recovery, in 2017 (Wisconsin Voices For Recovery,, 2021). This study addresses analytic questions that the MARI program could not directly study. For example, the MARI team only had data on MARI participants and could not address broader Dane County outcomes (White et al.,, 2021; Nyland et al.,, 2024). Additionally, since the program was not implemented in isolation, it is difficult to assess Dane County-level trends due to MARI or ED2Recovery. Therefore, the DES simulation model provides insight into the effects of the three different OUD treatment programs.
Table 1 reports model inputs used in the case study. Table 1 includes distributions, model inputs, data sources used, and parameter estimation type. The distributions include: exponential (Exp) with parameter $\lambda$, triangular (Tri) with parameters minimum, median, and maximum, and log-normal (LN) with parameters $\mu$, $\sigma$, a location parameter of 0 , and a truncation parameter of infinity unless otherwise stated. A LN distribution is used to estimate times in OUD treatment, CJS, and the hospital, since individuals are most likely to spend a short amount of time in jail, OUD treatment, and the hospital with a low likelihood of remaining in each of these states for an extended period of time. Due to the “new wave” of opioid deaths starting in 2019 caused by the introduction of fentanyl in the drug supply, we split arc (2) into pre-2019 (i.e., through the end of 2018) and post-2019 (i.e., 2019 and on) to capture the increased deadliness of opioids due to fentanyl (National Center for Health Statistics,, 2022).
@@
Dist. Parameter Distribution for Base Model Data Source Parameter
Est. Type
NA Age (years) at opioid initiation Truncated LN(2.08, 0.76) SAMHSA (2019) D
with Loc = 12 and Trunc=105
NA Age (years) in starting population Truncated LN(3.74, 0.49) SAMHSA (2019) D
with Loc = 12 and Trunc=105
NA Starting population size Tri(27299, 34224, 43261) SAMHSA (2019) D
NA Starting state Multinomial, See Section 2.4 Expert Opinion C
(1) Time of next arrival (days) Exp(10.87) Multiple years 2015-2019 SAMHSA (2019) D
@@
(2) pre-2019 Time in active state given next event is an opioid-related death before 2019 (days) Right-skewed non-parametric version of LN(17.60, 4.19) National Center for Health Statistics, (2022) 2016-2018 DM
@@
(2) post-2019 Time in active state given next event is an opioid-related death 2019 and later (days) Right-skewed non-parametric version of LN(17.13, 4.14) National Center for Health Statistics, (2022) 2019-2020 DM
@@
(3) Time in active state given next event is a hospital encounter (days) Right-skewed non-parametric version of LN(9.07, 3.01) Wisconsin Department of Health Services, 2017b 2018-2019 DM
@@
(4) Time in active state given next event is an opioid-related arrest (days) Right-skewed non-parametric version of LN(10.04, 2.35) Wisconsin Department of Justice, (2016) 2016-2019 DM
@@
(5) Time in active state given next event is OUD treatment (days) Right-skewed non-parametric version of LN(7.48, 1.03) Wisconsin Department of Health Services, (2019) 2016-2017 DM
@@
(6) Time in active state given next event is inactive use (days) Right-skewed non-parametric version of LN(4.82, 2.20) Rivera et al., (2018); Bauer, (2019) DM
@@
(7) Time until individual's non-opioid related death (days) Empirical, See Section A.5.7 Arias et al., (2013) D
@@
(8) Time until individual's non-opioid related arrest (days) Right-skewed non-parametric version of LN(7.88, 2.38) Wisconsin Department of Justice, (2016) 2016-2019 DM
(A) Hospital service time (days) LN(0.82, 0.48) Singh and Cleveland, (2020); Reinert et al., (2019) D
(B) CJS service time (days) LN(2.16, 1.47) Zhang et al., (2022) D
@@
(C) OUD treatment service time (days) LN(4.78, 1.18) Division of Care and Teatment Services, (2018) D
@@
(D) Inactive service time after release from CJS (days) LN(3.29, 1.61) Bukten et al., (2017); Kinlock et al., (2008) D
(E) Inactive service time after OUD treatment (days) LN(4.52, 1.09) Nunes et al., (2018) D
@@
(F) Inactive service time after hospital or ED discharge (days) LN(1.95, 1.40) Chutuape et al., (2001) D
@@
(G) Inactive service time after active state (days) LN(6.29, 2.51) Calibrated C
$p_{D}$ of Hosp. encounters leading to death Probability(0.0218) Singh and Cleveland, (2020) D
$p_{A}$ of Hosp. encounters leading to arrest Probability(0.01) Expert Opinion C
$p_{AD}$ of Arrests leading to treatment Probability(0) Not Applicable NA
$p_{OD}$ of Hosp. encounters leading to treatment Probability(0.2227) Singh and Cleveland, (2020) D
$p_{CM}$ of CJS re-entries leading to treatment Probability(0) Not Applicable NA
Loc: Location Parameter, Trunc: Truncation Parameter
Parameter Estimation Type has four possible types. D: Directly estimated. DM: Directly estimated with multiple estimation approaches considered. C: Calibrated using expert opinion as a starting point. NA: Not applicable.
@@
Table 2 reports the 2017 USD average cost per event for opioid-related-death, opioid-related-arrest, start OUD treatment, hospital encounter, active opioid use at year-end, and inactive opioid use at year-end. We obtain the 2017 USD cost of a single opioid-related death of \$11,548,462 from Luo et al., (2021). We estimate the cost of an individual starting OUD treatment in Wisconsin to be \$$\$8,224$. The per-event cost of starting OUD treatment was estimated by dividing the total estimated Wisconsin burden in 2017 due to OUD treatment (i.e., \$$\$271.4$ million) by the total number of 2017 OUD treatment episodes in the state of Wisconsin (i.e., $33,005$) obtained from Wisconsin Department of Health Services, (2019). The total estimated Wisconsin burden in 2017 due to OUD treatment was estimated by re-allocating the 2017 WI healthcare costs in Luo et al., (2021) to be 53.93% healthcare cost and 46.07% OUD treatment based on Table A.3 in SAMHSA (2019). All hospitals, physicians, and other professionals were attributed to hospital encounter costs. Free-standing nursing homes, free-standing home health, and other residential, personal, and public health costs were attributed to OUD treatment episodes. Retail prescription drug costs and insurance administration were split evenly between hospital encounters and OUD treatment costs. Therefore, we estimated the cost of a hospital encounter to be \$$\${12,051}$. Similarly, the per-event cost of a hospital encounter was estimated by dividing the total estimated Wisconsin burden in 2017 due to health care (i.e., \$$\$317.8$ million) by the total number of 2017 hospital encounters in the state of Wisconsin (i.e., $26,369$) from Wisconsin Department of Health Services, 2017b .
@@
Start OUD treatment \$8,224
Hospital encounter \$12,051